how do I solve this question
An arithmetic progression is a sequence of numbers whose nth term is given by a + (n - 1)d for some a and d. The sequence of numbers below form an arithmetic progression. 3, 7, 11, 15, 19, ... Which values of a and d lead to this sequence of numbers?
Select one:

A. a = 3 and d = 4
B. a = 4 and d = 3
C. a = 3 and d = 7
D. a = 7 and d = 3

Found 2 solutions by richwmiller, chndnbl:
Answer by richwmiller(17219)

the first value is 3 and the d is 4 ie add 4 to get next number
an = a1 + (n - 1)d
A. a = 3 and d = 4

Answer by chndnbl(31)

first term a=3
difference d = 7-3=11-7=15-11 so on =4
hence
a=3 and d=4